# Imaginaries, products and the adele ring

### v2

@@ 4. Imaginaries of atomless Boolean powers
($\clubsuit$). Assume T{T} admits EI. Then $T_{atomless}$ admits weak EI.
@@
We may assume by Morleyzation that $T$ admits QE. Let $M$ be a sufficiently saturated and homogeneous model of $T_{atomless}$.
Construct $N\models T_{at}$ as follows: Let $\widehat{B}$ be the Stone space of $B(M)$. For $p\in\widehat{B}$, let $M_{p}$ be the quotient by the maximal ideal corresponding to $p$; this is a 2-valued model of $T$. Let $N=\prod_{p\in\widehat{B}}M_{p}$. In particular $B(N)$ is the Boolean algebra of all subsets of $\widehat{B}$. Then $N\models T_{at}$ (Example 2.8). We have a natural embedding of $M$ into $N$, mapping $m$ to $p\mapsto m_{p}$ where $m_{p}$ is the image of $m$ in $M_{p}$.
Define $\widetilde{N}$}} to be $N/J$, where $J$ is the atomic ideal of $\widehat{B}$. Then ${{\widetilde{N}}}\models T_{atomless}$. We have a natural map $N\to{{\widetilde{N}}}$; let $\mathbf{n}$} denote the image of $n$ in $\widetilde{N}$}}. Let $\alpha:M\to{{\widetilde{N}}}$ be the composition, and denote $\alpha(m)$ by $\mathbf{m}$}. Then $\alpha$ is injective: if $m,m^{\prime}$ agree away from finitely many ultrafilters of $B$, then $m=m^{\prime}$, since otherwise $m\neq m^{\prime}]$ is a nonzero element of $B$ and as $B$ is atomless, it lies in infinitely many ultrafilters. For the same reason it is an $L^{boole}$-embedding, and hence elementary.
@@
Consider first the case where $n=b\in B(N)$. The formula responsible for the algebraicity of $b$ over $m$ can be taken to have the form $\Psi(y,[\phi_{1}(m)],\cdots,[\phi_{l}(m)])$, with $\Psi$ a formula in the language of Boolean algebras, by the quantifier-elimination of Theorem 2.3. Thus in this case it suffices to show that within $B(N)$, if $b\in\mathop{\rm acl}\nolimits_{N}(c_{1},\ldots,c_{l})$ then ${\mathbf{b}}\in\mathop{\rm dcl}\nolimits({\mathbf{c}}_{1}),\ldots,{\mathbf{c}}_{l}))$. This is easy to see explicitly; the algebraic closure within $B(N)$ of $c_{1},\ldots,c_{l}$ is the smallest Boolean subalgebra $C$ of $B(N)$ including $c_{1},\ldots,c_{l}$ and such that if a sum of finitely many atoms $a_{1},\ldots,a_{\mu}$ of $B(N)$ lies in $C$, then so does each $a_{i}$. The Boolean operators reduce homomorphically to $B/J$, while the operation on atoms does not increase $A$ modulo $J$ since each $a_{i}$ lies in $J$.
Now allow $n$ to lie in a product sort of $N$, and consider the formula responsible for the algebraicity of $n$ over $m$; it can be taken to have the form $\Theta([\phi_{1}(m,x)],\cdots,[\phi_{l}(m,x)])$ by Theorem 2.3; where $\Theta$ is a Boolean algebra formula, and $\phi_{1},\ldots,\phi_{l}$ are formulas of $L$. Let $b_{i}=[\phi_{i}(m,x)]$. Then $b_{i}\in\mathop{\rm acl}\nolimits_{N}(m)$, so $\alpha(b_{i})\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$ by the previous case. We can take $\phi_{1},\ldots,\phi_{l}$ to be a partition of $1$, so that the $b_{i}$ also form a partition of $1$ (but some of them may be zero.) Now there are only finitely many solutions of $\bigwedge_{i\leq l}[\phi_{i}(m,x)]=b_{i}$. It follows that for for some $k_{i}$ we have $b_{i}\leq[(\exists^{\leq k_{i}}x)\phi_{i}(m,x)]$. Replacing the $b_{i}$ by a finer partition according to the various formulas $\exists^{\leq l}x$, $l\leq k_{i}$, we may assume $b_{i}\leq[(\exists^{!k_{i}}x)\phi_{i}(m,x)]$. Moreover for each $i$, either $b_{i}=0$ or $b_{i}$ is a finite union of atoms, or else $k_{i}=1$. (Otherwise we can find infinitely many functions on $\widehat{B}$ satisfying $\phi_{i}(x,m)]=b_{i}$.) Now moving to $\widetilde{N}$}}, the $\phi_{i}(m,n)]$ where $k_{i}>1$ vanish, and it becomes clear that ${\mathbf{n}}\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}},\alpha(b_{1}),\ldots,\alpha(b_{l}))$. Since $b_{i}\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$, we are done. ∎
@@
Let $n\in N$, and assume $n\in\mathop{\rm acl}\nolimits_{N}(m)$ for some tuple $m$ from $M$. Then ${\mathbf{n}}\in\mathop{\rm acl}\nolimits_{\hat{N}}({\mathbf{m}})$ by Claim 1. Since $\alpha(M)\prec\hat{N}$ is elementary, $\mathbf{n}$} is the image of an element $n^{\prime}\in M$. Then $\alpha(n)={\mathbf{n}}=\alpha(n^{\prime})$ so $n,n^{\prime}$ differ only at finitely many elements of $\widehat{B}$. This finite subset of $\widehat{B}$ is thus $m$-invariant; by stable embeddedness of $B$, we have a finite set of ultrafilters invariant under $Aut(B/b)$ for some finite tuple $b$ from $B$; this is easily seen to be impossible unless the set is empty.
Turning now to the statement of Theorem 4.1, let $E$ be a 0-definable equivalence relation on a 0-definable set $D$. By quantifier-elimination for $T_{atomless}$ Theorem 2.3, we may assume $E$ is defined by a quantifier-free formula. Then $E$ continues to define an equivalence relation on any substructure of a model of $T_{atomless}$.
Let $d\in D(M)$. We wish to code the $E$-class $\bar{d}$} of $d$.
Using Theorem 3.1, let $e$ be a finite subset of $N$ such that $d^{\prime}\in\mathop{\rm dcl}\nolimits_{N}(e)$ and $e\subset acl_{N}(d^{\prime})$; where $d^{\prime}$ is $d$ viewed as an element of $N$. By Claim 2, we have $e\subset M$. Now any automorphism of $M$ extends to $Aut(N)$; hence $e\subset acl_{M}({\bar{d}})$, and ${\bar{d}}\in\mathop{\rm dcl}\nolimits_{M}(e)$. ∎
4.2. Variations on the proof
It would also be possible to avoid going through $\widetilde{N}$}, and deduce Claim 2 directly from the ‘Moreover’ of the proof of Claim 1; arguing instead that the universal parts of $T_{at}$ and of $T_{atomless}$ coincide. Going further in this direction, it would be possible to verify the hypothesis of the following remark, adapted to weak EI and algebraic closure.
@@ 4.2. Variations on the proof
Let $T$ be a theory, whose universal part $T_{\forall}$ admits a model companion $\widetilde{T}$. Assume $T$ admits quantifier-free-EI in the following sense: if $D$ is a quantifier-free definable set and $E$ a quantifier-free definable equivalence relation on $D$ (with no parameters), then there exists in $T$ a definable function $f$ on $D$ such $f(x)=u$ is quantifier-free definable, and $xEy$ if and only if $f(x)=f(y)$.
Then $\widetilde{T}$ admits EI.
@@
Consider a quantifier-free definable equivalence relation $E$ on a quantifier-free definable set $D$, for the theory $\widetilde{T}$. Then $E$ is still an equivalence relation for $T$, since the universal parts of $\widetilde{T}$ and of $T$ are the same. Now in $T$ we have a definable function $f$ on $D$ such $f(x)=u$ is quantifier-free definable, and $xEy$ iff $f(x)=f(y)$. So $f$ defines a function in models of $\widetilde{T}$ too (as they are existentially closed $f$ is total on $D$), and the same equation holds. ∎
@@
Fix an $L$-theory $T$. Recall the definition of $L_{mixed}$ (§ 2.6(3).)
@@
Construct $E^{Z}$ as in § 3.15; see Remark 3.16. Then Lemma 3.17 continues to hold: $Z$ is a union of $E^{Z}$-classes, i.e. for $x,y\in c$, $xE^{Z}y$ implies $x\in Z\iff y\in Z$. This is easy to see in models where $J$ is the ideal generated by the atoms of $B$, by induction on the number of atoms in $x\triangle y$. Since such models are elementarily dense in the class of models of $T_{mixed}$ (Lemma 2.9 (2)), it holds in general.
@@
Let $M^{[n]}$ denote the set of $n$-element subsets of $M$ (or the set of distinct $n$-tuples modulo $Sym(n)$.).
@@
By full embeddedness of $M/J$ (Lemma 2.7), the equivalence relation $f(x)=f(y)$ on $X$ is $M/J$-definable (with parameters.) By factoring out this equivalence relation, we may assume $f$ is 1-1. Using weak elimination of imaginaries for $T_{atomless}$ (Theorem 4.1) applied to $M/J$, we may assume $X\subset(D/J)^{[m]}$ for some definable set $D$ of $L$. Note $(D/J)^{m}$ can be identified with $D^{m}/J$; the identification is compatible with the $Sym(m)$-actions. Pulling $f$ back to $(D/J)^{m}=D^{m}/J$, we find a definable function $F$ with domain ${\mathop{\operatorname{\rm dom}}}(F)\subset D^{m}$ and range contained in $M^{[n]}$, such that $F(x)=F(y)$ iff $xJy^{\sigma}$ for some $\sigma\in Sym(m)$.
@@
By Lemma 2.9(4), $(\exists^{\infty})$ is definable in $T_{atomless}$. Hence if the domain can be infinite in some model, then an example of this exists in any model. Thus we may assume $M=(\prod_{i\in I}M_{i},{\mathop{\operatorname{\rm P}}}(I),J)$ with $I$ and the $M_{i}$ countable, and $J$ the ideal generated by the atoms. Note $Sym(m)$ respects $J$ on $D^{m}(M)$, i.e. the two equivalence relations commute: if $x^{\sigma}Jy$ then $xJy^{\sigma^{-1}}$. We give $M$ the natural Borel structure (e.g. taking $M_{i}$ to have universe $\subset{\mathbb{N}}$ and thus embedding $M$ in $\mathbb{N}$$^{\mathbb{N}}$.) We remark also that ${\mathop{\operatorname{\rm dom}}}(F)$ is Borel by QE in $T_{atomless}$ (though we will not really need this fact, Borelness of the ambient $D(M)^{m}$ suffices.)
@@ 5.4. Adeles
Conversely, $B$ is interpretable as the set of idempotent finite adeles, and by [8, Theorem 5.2] the set $I_{fin}\subset B$ is definable from the ring structure on ${\mathbb{A}}^{fin}_{{\mathbb{Q}}}$. Note also that $\prod_{p}{\mathbb{Q}}_{p}$ is interpretable in $\prod_{p}{\mathbb{Z}}_{p}$ as the localization by the multiplicative set of all non-zero divisors.
@@
Each ${\mathbb{Q}}_{p}$ eliminates imaginaries in the language with the field sort, and a sort $S_{n}$ interpreted as the set of lattices, $GL_{n}(K)/GL_{n}({\mathcal{O}})$ (see [17]). We require however uniform elimination of imaginaries over the class of all ${\mathbb{Q}}_{p}$; this amounts to EI for each ${\mathbb{Q}}_{p}$ separately, as well as for each nonprincipal ultrproduct of the ${\mathbb{Q}}_{p}$. The latter is achieved in [17] only over certain field constants; this is greatly ameliorated in [11], where only algebraic constants are required. In order to make use of this, we now expand the language in two ways.
@@
Moreover, the Galois group of the degree $n$ extension of the residue field is uniformly interpretable over $k$; we add a constant symbol $\mathfrak{f}_{n}$ to this finite sort $\mathcal{Gal}_{n}$, and interpret it in ${\mathbb{Q}}_{p}$ as the class of the Frobenius automorphism. See [4], [14].
@@
View ${\mathbb{Q}}_{p}$ as a $PL$-structure; then the constants are naturally interpreted in ${\mathbb{A}}^{fin}_{{\mathbb{Q}}}$. In particular $\tau_{n}$ is interpreted by the image of element $(2,3,5,\cdots)$.
@@
It is shown in [11] (Lemma 6.1.7) that the ${\mathbb{Q}}_{p}$ uniformly weakly eliminate imaginaries in this language. From this and Theorem 5.1 the result follows as in the proof of Corollary 5.5. ∎
@@
If $u$ is a finite 0-definable set, and $X$ another 0-definable set, we can consider the twisted power $X^{u}$. $X^{u}$ is the set of $u$-tuples from $X$, i.e. the set of functions from $u$ to $X$. It is a form of the cartesian power $X^{|u|}$, in the sense that $X^{u}$ becomes definably isomorphic to $X^{|u|}$ once the elements of $u$ are named. (See [16] for example.)
@@
Assume the expansion by constants for elements of $\cup_{u\in\mathcal{U}}u$ admits elimination of imaginaries. Then $T$ admits elimination of imaginaries to the sorts $X^{u}$ ($X\in\mathcal{X},\ u\in\mathcal{U}$).
@@
This is a consequence of [15, Lemma 2.5], but for convenience we repeat the proof. Let $e$ be an imaginary. By the elimination of imaginaries assumption, there exists $u\in\mathcal{U}$, and $c\in u$ such that $dcl(e,c)=dcl(d,c)$ with $d\in X=\prod_{i=1}^{m}X_{i}$, $X_{i}\in\mathcal{X}$.

### v3

@@ 4. Imaginaries of atomless Boolean-valued models
($\clubsuit$). Assume T{T} admits EI. Then $T_{atomless}^{boole}$ admits weak EI.
@@
We may assume by Morleyzation that $T$ admits QE. Let $M$ be a sufficiently saturated and homogeneous model of $T_{atomless}^{boole}$.
Construct $N\models T_{at}$ as follows: Let $\widehat{B}$ be the Stone space of $B(M)$. For $p\in\widehat{B}$, let $M_{p}$ be the quotient by the maximal ideal corresponding to $p$; this is a 2-valued model of $T$. Let $N=\prod_{p\in\widehat{B}}M_{p}$. In particular $B(N)$ is the Boolean algebra of all subsets of $\widehat{B}$. Then $N\models T_{at}$ (see Example 2.9). We have a natural embedding of $M$ into $N$, mapping $m$ to $p\mapsto m_{p}$ where $m_{p}$ is the image of $m$ in $M_{p}$.
Define $\widetilde{N}$}} to be $N/J$, where $J$ is the atomic ideal of $\widehat{B}$. Then ${{\widetilde{N}}}\models T_{atomless}^{boole}$ (cf. Lemma 2.8). We have a natural map $N\to{{\widetilde{N}}}$; let $\mathbf{n}$} denote the image of $n$ in $\widetilde{N}$}}. Let $\alpha:M\to{{\widetilde{N}}}$ be the composition, and denote $\alpha(m)$ by $\mathbf{m}$}. Then $\alpha$ is injective: if $m,m^{\prime}$ agree away from finitely many ultrafilters of $B$, then $m=m^{\prime}$, since otherwise $m\neq m^{\prime}]$ is a nonzero element of $B$ and as $B$ is atomless, it lies in infinitely many ultrafilters. For the same reason it is an $L^{boole}$-embedding, and hence elementary.
@@
Consider first the case where $n=b\in B(N)$. By the quantifier-elimination of Theorem 2.3 the formula responsible for the algebraicity of $b$ over $m$ can be taken to have the form $\Psi(y,[\phi_{1}(m)],\cdots,[\phi_{l}(m)])$, where $\phi_{i}$ are $L$-formulas and $\Psi$ is a formula in the language of Boolean algebras. Thus in this case it suffices to show that within $B(N)$, if $b\in\mathop{\rm acl}\nolimits_{N}(c_{1},\ldots,c_{l})$ then ${\mathbf{b}}\in\mathop{\rm dcl}\nolimits({\mathbf{c}}_{1},\ldots,{\mathbf{c}}_{l}))$. One can verify this explicitly; the algebraic closure within $B(N)$ of $c_{1},\ldots,c_{l}$ is the smallest Boolean subalgebra $A$ of $B(N)$ including $c_{1},\ldots,c_{l}$ and such that if a sum of finitely many atoms $a_{1},\ldots,a_{\mu}$ of $B(N)$ lies in $A$, then so does each $a_{i}$. The Boolean operators reduce homomorphically to $B/J$, while the operation on atoms does not increase $A$ modulo $J$ since each $a_{i}$ lies in $J$.
Now allow $n$ to lie in a product sort of $N$, and consider the formula responsible for the algebraicity of $n$ over $m$; it can be taken to have the form $\Theta([\phi_{1}(m,x)],\cdots,[\phi_{l}(m,x)])$ by Theorem 2.3; where $\Theta$ is a Boolean algebra formula, and $\phi_{1},\ldots,\phi_{l}$ are formulas of $L$. Let $b_{i}=[\phi_{i}(m,n)]$. As $n\in\mathop{\rm acl}\nolimits(m)$, we have $b_{i}\in\mathop{\rm acl}\nolimits_{N}(m,n)=\mathop{\rm acl}\nolimits_{N}(m)$, so $\alpha(b_{i})\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$ by the previous case. We can take $\phi_{1},\ldots,\phi_{l}$ to be a partition of $1$, so that the $b_{i}$ also form a partition of $1$ (but some of them may be zero.)
There are only finitely many solutions of $\bigwedge_{i\leq l}[\phi_{i}(m,x)]=b_{i}$ since all of them satisfy $\Theta([\phi_{1}(m,x)],\cdots,[\phi_{l}(m,x)])$. It follows that for some $k_{i}$ we have $b_{i}\leq[(\exists^{\leq k_{i}}x)\phi_{i}(m,x)]$. Replacing the $b_{i}$ by a finer partition according to the various formulas $\exists^{\leq l}x$, $l\leq k_{i}$, we may assume $b_{i}\leq[(\exists^{!k_{i}}x)\phi_{i}(m,x)]$. Moreover for each $i$, either $b_{i}=0$ or $b_{i}$ is a finite union of atoms, or else $k_{i}=1$. (Otherwise we can find infinitely many functions on $\widehat{B}$ satisfying $\phi_{i}(x,m)]=b_{i}$). Now moving to $\widetilde{N}$}}, the $\phi_{i}(m,n)]$ where $k_{i}>1$ vanish, and it becomes clear that ${\mathbf{n}}\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}},{\mathbf{b}}_{1},\ldots,{\mathbf{b}}_{l})$. Since $b_{i}\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$, we have ${\mathbf{n}}\in\mathop{\rm dcl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$. ∎
@@
Let $n\in N$, and assume $n\in\mathop{\rm acl}\nolimits_{N}(m)$ for some tuple $m$ from $M$. Then ${\mathbf{n}}\in\mathop{\rm acl}\nolimits_{{{\widetilde{N}}}}({\mathbf{m}})$ by Claim 1. Since $\alpha(M)\prec{{\widetilde{N}}}$ is elementary, $\mathbf{n}$} is the image of an element $n^{\prime}\in M$. Then $\alpha(n)={\mathbf{n}}=\alpha(n^{\prime})$ so $n,n^{\prime}$ differ only at finitely many elements of $\widehat{B}$. This finite subset of $\widehat{B}$ is thus $m$-invariant; by stable embeddedness of $B$, (cf. Theorem 2.3), we have a finite set of ultrafilters invariant under $Aut(B/b)$ for some finite tuple $b$ from $B$; this is easily seen to be impossible unless the set is empty. ∎
Claim 3.   $M,N$ have the same universal theory.
Proof.
[…]
Turning now to the statement of Theorem 4.1, let $E$ be a 0-definable equivalence relation on a 0-definable set $D$. By quantifier-elimination for $T_{atomless}^{boole}$ (Theorem 2.3 combined with quantifier elimination for the Boolean sort $B$ given in Lemma 2.5(1)), we may assume $E$ is defined by a quantifier-free formula. Then $E$ continues to define an equivalence relation on any model of the universal theory of $M$, in particular on $N$ by Claim 3.
Let $d\in D(M)$. We wish to code the $E$-class $\bar{d}$} of $d$. Let $d^{\prime}\in N^{eq}$ code the $E$-class of $d$ in $N$.
Using Theorem 3.1, let $e$ be a finite subset of $N$ such that $d^{\prime}\in\mathop{\rm dcl}\nolimits_{N}(e)$ and $e\subset acl_{N}(d^{\prime})$. By Claim 2, we have $e\subset M$. Now any automorphism $\tau$ of $M$ extends to $\tau^{\prime}\in Aut(N)$; thus if $\tau(e)=e$ then $\tau^{\prime}(d^{\prime})=d^{\prime}$ so $\tau(d)Ed$; hence ${\bar{d}}\in\mathop{\rm dcl}\nolimits_{M}(e)$. Similarly $e$ cannot have infinitely many distinct $Aut(M/{\bar{d}})$-conjugates, so $e\subset acl_{M}({\bar{d}})$. This proves weak EI in $M$. ∎
@@
It would also be possible to avoid Claim 1 and going through $\widetilde{N}$}, and deduce Claim 2 directly along the lines of Claim 3 and the proof of the ‘Moreover’ towards the end of the proof of Claim 1.
Say $T$ admits quantifier-free-EI if whenever $D$ is a quantifier-free definable set and $E$ a quantifier-free definable equivalence relation on $D$ (with no parameters), then there exists in $T$ a quantifier-free partition $D=\bigcup_{i=1}^{k}D_{i}$, and definable function $f_{i}$ on $D_{i}$ such $f_{i}(x)=u$ is quantifier-free definable, and for $x,y\in D_{i}$, $xEy$ if and only if $f(x)=f(y)$.
Remark 4.3.
Let $T$ be a theory, whose universal part $T_{\forall}$ admits a model completion $\widetilde{T}$. Assume $T$ admits quantifier-free-EI. Then $\widetilde{T}$ admits EI.
@@
Consider a quantifier-free definable equivalence relation $E$ on a quantifier-free definable set $D$, for the theory $\widetilde{T}$. Then $E$ is still an equivalence relation for $T$, since the universal parts of $\widetilde{T}$ and of $T$ are the same. Now in $T$ we have a definable function $f$ on $D$ such $f(x)=u$ is quantifier-free definable, and $xEy$ iff $f(x)=f(y)$. So $f$ defines a function in models $\tilde{N}$ of $\widetilde{T}$ (as $\tilde{N}$ is existentially closed, and extends to a model of $T$, $f_{i}$ has domain $D_{i}$ in $\tilde{N}$), and the same equation holds. ∎
@@
Fix an $L$-theory $T$. Recall the definition of $L_{mixed}$ in § 2.6(3).
@@
Construct $E^{Z}$ as in § 3.15; see Remark 3.16. Then Lemma 3.17 continues to hold: $Z$ is a union of $E^{Z}$-classes, i.e. for any $x,y$ we have $xE^{Z}y$ implies $x\in Z\iff y\in Z$. This is easy to see in models where $J$ is the ideal generated by the atoms of $B$, by induction on the number of atoms in $x\triangle y$. Since such models are elementarily dense in the class of models of $T_{mixed}$ (Lemma 2.10 (2)), it holds in general.
@@
Let $M^{[n]}$ denote the set of $n$-element subsets of $M$ (or the set of distinct $n$-tuples modulo $Sym(n)$).
@@
By full embeddedness of $M/J$ (Lemma 2.8), the equivalence relation $f(x)=f(y)$ on $X$ is $M/J$-definable (with parameters.) By factoring out this equivalence relation, we may assume $f$ is 1-1. Using weak elimination of imaginaries for $T_{atomless}^{boole}$ (Theorem 4.1) applied to $M/J$, we may assume $X\subset(D/J)^{[m]}$ for some definable set $D$ of $L$. Note $(D/J)^{m}$ can be identified with $D^{m}/J$; the identification is compatible with the $Sym(m)$-actions. Pulling $f$ back to $(D/J)^{m}=D^{m}/J$, we find a definable function $F$ with domain ${\mathop{\operatorname{\rm dom}}}(F)\subset D^{m}$ and range contained in $M^{[n]}$, such that $F(x)=F(y)$ iff $xJy^{\sigma}$ for some $\sigma\in Sym(m)$.
@@
By Lemma 2.10(4), $(\exists^{\infty})$ is definable in $T_{atomless}^{boole}$. Hence if the domain can be infinite in some model, then an example of this exists in any model. Thus we may assume $M=(\prod_{i\in I}M_{i},{\mathop{\operatorname{\rm P}}}(I),J)$ with $I$ and the $M_{i}$ countable, and $J$ the ideal generated by the atoms. Note $Sym(m)$ respects $J$ on $D^{m}(M)$, i.e. the two equivalence relations commute: if $x^{\sigma}Jy$ then $xJy^{\sigma^{-1}}$. We give $M$ the natural Borel structure (e.g. taking $M_{i}$ to have universe $\subset{\mathbb{N}}$ and thus embedding $M$ in $\mathbb{N}$$^{\mathbb{N}}$). We remark also that ${\mathop{\operatorname{\rm dom}}}(F)$ is Borel by QE in $T_{atomless}^{boole}$ (though we will not really need this fact, Borelness of the ambient $D(M)^{m}$ suffices).
@@ 5.4. Adeles
Conversely, $B$ is interpretable as the set of idempotent finite adeles, and by [9, Theorem 5.2] the set $I_{fin}\subset B$ is definable from the ring structure on ${\mathbb{A}}^{fin}_{{\mathbb{Q}}}$. Note also that $\prod_{p}{\mathbb{Q}}_{p}$ is interpretable in $\prod_{p}{\mathbb{Z}}_{p}$ as the localization by the multiplicative set of all non-zero divisors.
@@
Each ${\mathbb{Q}}_{p}$ eliminates imaginaries in the language with the field sort, and a sort $S_{n}$ interpreted as the set of lattices, $GL_{n}(K)/GL_{n}({\mathcal{O}})$ (see [18]). We require however uniform elimination of imaginaries over the class of all ${\mathbb{Q}}_{p}$; this amounts to EI for each ${\mathbb{Q}}_{p}$ separately, as well as for each nonprincipal ultraproduct of the ${\mathbb{Q}}_{p}$. The latter is achieved in [18] only over certain field constants; this is greatly ameliorated in [12], where only algebraic constants are required. In order to make use of this, we now expand the language in two ways.
@@
Moreover, the Galois group of the degree $n$ extension of the residue field is uniformly interpretable over $k$; we add a constant symbol $\mathfrak{f}_{n}$ to this finite sort $\mathcal{Gal}_{n}$, and interpret it in ${\mathbb{Q}}_{p}$ as the class of the Frobenius automorphism. See [5] and [15].
@@
View ${\mathbb{Q}}_{p}$ as a $PL$-structure; then the constants are naturally interpreted in ${\mathbb{A}}^{fin}_{{\mathbb{Q}}}$. In particular $\tau_{n}$ is interpreted by the image of the element $(2,3,5,\cdots)$.
@@
It is shown in [12, Cor. 6.1.7] that ${\mathbb{Q}}_{p}$ uniformly eliminate imaginaries in this language. From this and Theorem 5.1 the result follows as in the proof of Corollary 5.5. ∎
@@
If $u$ is a finite 0-definable set, and $X$ another 0-definable set, we can consider the twisted power $X^{u}$.
Definition 5.7.
$X^{u}$ is the set of $u$-tuples from $X$, i.e. the set of functions from $u$ to $X$.
$X^{u}$ is a form of the cartesian power $X^{|u|}$, in the sense that $X^{u}$ becomes definably isomorphic to $X^{|u|}$ once the elements of $u$ are named. (See [17] for example.)
@@
Assume the expansion by constants for elements of $\bigcup_{u\in\mathcal{U}}u$ admits elimination of imaginaries. Then $T$ admits elimination of imaginaries to the sorts $X^{u}$ (with $X\in\mathcal{X},\ u\in\mathcal{U}$).
@@
This is a consequence of [16, Lemma 2.5], but for convenience we repeat the proof. Let $e$ be an imaginary. By the elimination of imaginaries assumption, there exists $u\in\mathcal{U}$, and $c\in u$ such that $dcl(e,c)=dcl(d,c)$ with $d\in X=\prod_{i=1}^{m}X_{i}$, $X_{i}\in\mathcal{X}$.